Kay started a computer club, and for a while she was the only member. She planned to have each member find two new members each month. By the end of the first month she had found two new members. If her plan is carried out, how many members will the club have at the end of 10 months



1
Expert's answer
2021-02-03T03:58:12-0500

At the begiining: "n_0=1=3^0".

At the end of the 1st month : "n_1=3=3^1".

At the end of the 2nd month : "n_2=9=3^2".

"..."

At the end of the 10th month : "n_{10}=3^{10}=59049".


The club will have 59049 members at the end of 10 months.